OpenStudy (anonymous):

Solve the system of equations: 5x-3y=-34 2x+7y=-30

OpenStudy (anonymous):

Have you learnt the elimination method?

hero (hero):

Multiply the first equation by 2: 10x - 6y = -68 2x + 7y = -30 Combine both equations together: 12x + y = -98 Isolate y: y = -98 - 12x substitute the expression for y into the second equation: 2x + 7(-98 - 12x) = -30 2x -686 - 84x = -30 -686 + 30 = 84x - 2x -656 = 82x -656/82 = x -8 = x

hero (hero):

Now you should be able to find y on your own by substituting x = -8 back into one of the original equations.

hero (hero):

@dcortinas, have you figured out the value of y yet?